A block weighing 130N is an incline,whose slope is 5 vertical to horizontal. It`s initial velocity down the incline is 2.4 m/S . What will be the velocity 5 seconds later ? Take coefficient of friction at contact surface=0.3

Grade:12th pass

1 Answers

Gauthamkrishna Nemani
7 years ago
The forces acting on the body that determines the velocity of the body will be mgsinx downward along the incline and frictional force along the surface upwardsassuming vertical to horizontal means tanxTanx=5Sinx=5/√24so ,ma= mgsinx- umgcosx. (u is the coefficient of friction and mgcosx is the normal reaction as the force acting normal to the inclined surface)Given mg=130 ie m=13kg sinx=5/√24 cosx=1/√24On substituting the values we find `a` and to find velocity we use v=u+at